OpenStudy (anonymous):

Tickets numbered 1 to 20 are mixed up and then a ticket is drawn at random. What is the probability that the ticket drawn has a number which is a multiple of 3 or 5? A. 1/2 B. 2/5 C. 8/15 D. 9/20

OpenStudy (anonymous):

d

OpenStudy (anonymous):

Here, S = {1, 2, 3, 4, ...., 19, 20}. Let E = event of getting a multiple of 3 or 5 = {3, 6 , 9, 12, 15, 18, 5, 10, 20}.

OpenStudy (anonymous):

well so.. Here, S = {1, 2, 3, 4, ...., 19, 20}. Let E = event of getting a multiple of 3 or 5 = {3, 6 , 9, 12, 15, 18, 5, 10, 20}. P(E) = n(E) / n(S) = 9 / 20 . Is this the answer?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank You... Have a good Time bro...

OpenStudy (anonymous):

u 2 man